If the sides of a square are lengthened by 5 cm, the area becomes 144cm^2. Find the length of a side of the original square.
The length of a side of the original square is____cm.

Answer by mananth(16946)

let original length be x
increased by 5
length = x+5
Are = S^2
(x+5)^2= 144
x^2+10x+25=144
x^2+10x-119=0
x^2+17x-7x-119=0
x(x+17)-7(x+17)=0
(x+17)(x-7)=0
x= 7 OR -17
Ignore negative value
x= 7 the original length